# ANIMALBOOTH: MULTIMODAL FEATURE ENHANCEMENT FOR ANIMAL SUBJECT PERSONALIZATION

### v1

@@ Abstract
Personalized animal image generation is challenging due to rich appearance cues and large morphological variability. Existing approaches often exhibit feature misalignment across domains, which leads to identity drift. We present AnimalBooth, a framework that strengthens identity preservation with an Animal Net and an adaptive attention module, mitigating cross domain alignment errors. We further introduce a frequency controlled feature integration module that applies Discrete Cosine Transform filtering in the latent space to guide the diffusion process, enabling a coarse to fine progression from global structure to detailed texture. To advance research in this area, we curate AnimalBench, a high resolution dataset for animal personalization. Extensive experiments show that AnimalBooth consistently outperforms strong baselines on multiple benchmarks and improves both identity fidelity and perceptual quality.
Index Terms—  personalized generation, animal subjects, diffusion models, ControlNet, generative AI.
@@ 1 INTRODUCTION
Personalized multimodal generation is a prominent yet challenging subfield that aims to synthesize images conforming to both textual descriptions (text–image consistency) and the intrinsic characteristics of custom concepts (identity consistency) [1, 2, 3, 4]. This paradigm shows strong potential across diverse applications ranging from creative artistry to product design [5, 6]. However, accurately capturing the identity of animal subjects, which exhibit complex visual features such as fine-grained fur textures and non-rigid morphologies, remains a significant hurdle [7, 8].
Fig. 1: Examples of Animal-HD dataset. Each instance provides five components: (a) a high-quality source image, (b) a detailed caption describing the scene and subject, (c) a precise pixel-level semantic segmentation mask of the primary subject, (d) the resulting masked subject for visual verification, and (e) the corresponding paired image.
@@
Existing personalization methodologies can be broadly categorized by their feature extraction and optimization strategies. The first category comprises optimization-based methods that achieve high fidelity by fine-tuning model components. A representative example is DreamBooth [9], which fine-tunes the UNet backbone, while Textual Inversion [10] instead optimizes word embeddings. This paradigm has been extended to support multiple customized concepts [11, 12, 13], including applications such as virtual dressing and editing [14]. Despite their fidelity, such approaches demand substantial computational resources and often exhibit an inherent trade-off between fidelity and diversity due to overfitting. The second category, encoder-based and tuning-free methods, offers a more efficient alternative. For instance, IP-Adapter [15] and its successors [16] utilize frozen external encoders (e.g., CLIP) to inject identity features. While efficient, their performance is fundamentally limited by the representational capacity of the encoder, particularly when modeling animals with unique identity cues such as body structures and fur patterns. Consequently, identity distortion and detail loss are frequent. Other directions include retrieval-augmented generation [17] or LLM-enhanced frameworks [18, 19, 20], yet the challenge of inadequate domain-specific representation remains.
@@
Motivated by bridging this domain gap, we propose AnimalBooth, a feature-enhanced, fine-tuning-free personalized generation framework specifically tailored for animals. Our approach adopts a dual-enhancement strategy: at the macroscopic level, we employ an Animal-Net with adaptive attention to reinforce structural and identity understanding; at the microscopic level, we introduce a frequency-controlled feature integration module, where Discrete Cosine Transform (DCT) filtering guides the diffusion process to preserve fine textures such as fur. Furthermore, while numerous datasets exist for general consistency, a dedicated benchmark for animal personalization is lacking. To address this, we construct AnimalBench, a high-resolution dataset for animal personalization, which will be released upon acceptance. Our contributions are threefold: (1) We propose a dedicated branch architecture comprising a lightweight Animal-Net and adaptive attention, effectively mitigating identity distortion. (2) We design a frequency-controlled feature integration module leveraging DCT filtering in latent space to enhance attribute manipulation and texture fidelity. (3) We establish and release AnimalBench, a high-definition dataset for animal personalization, on which our method achieves state-of-the-art performance.
Fig. 2: Qualitative comparison with other SOTA methods.
@@
Fig. 3: AnimalBooth primarily consists of a trainable Animal-Net and a frozen Photography-Net. The Animal-Net incorporates an Adaptive Attention module for efficient identity feature injection and a Frequency-Controlled module for enhanced control over visual attributes like structure and texture, while the Photography-Net integrates these features with text prompts within the latent space.
@@ 2.1 Overall Architecture
As depicted in Fig. 3, AnimalBooth effectively integrates a trainable Animal-Net with a frozen Photography-Net. The Animal-Net directly refines the complex features and flexible morphologies of animals within the feature space, effectively bypassing the complexities of cross-domain alignment. Through adaptive attention, identity features are efficiently injected. The incorporation of a frequency-controlled module further enhances fine-grained control over visual attributes such as the structure and textural details of the generated images. The frozen Photography-Net then merges these extracted fine-grained features with text prompts in the latent space.
@@ 2.2 Animal-Net
In the task of personalized animal image generation, precise extraction of fine-grained animal identity features is paramount for maintaining consistency between generated images and reference animals. To achieve this, we propose a specialized Animal-Net capable of simultaneously capturing both semantic information and high-frequency textural features of animals. Specifically, given an animal reference image $\mathcal{X}_{a}\in\mathbb{R}^{3\times H\times W}$, we first transform it into a latent space representation $\mathbf{L}_{a}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ using a frozen VAE encoder [21]. We then extract token embeddings from $\mathcal{X}_{a}$ using a frozen CLIP image encoder [22] and a trainable projection layer. Here, we employ Q-Former [23] as the projection layer. Subsequently, animal features within the Animal-Net interact extensively through a cross-attention mechanism, similar to the interaction between text and image features in the original Text-to-Image (T2I) model [21]. This interaction ensures a deep fusion of semantic and textural features. Finally, the output of the Animal-Net is aligned in parallel with the Photography-Net, injecting these fine-grained animal identity features into the Photography-Net via an adaptive attention module. Notably, the Animal-Net is solely used for encoding reference images; therefore, during the diffusion process, no noise is added to the reference image, and it undergoes only a single forward pass.
@@ 2.3 Adaptive Attention Module
To enable personalized animal image generation, the Photography-Net must possess both its original generative capabilities and the ability to fuse animal identity features. We freeze the core modules of the Photography-Net to preserve the former, and achieve the latter through an adaptive attention module. The architecture of the Photography-Net in AnimalBooth build upon SD v1.5 [21], with all self-attention modules replaced by adaptive attention modules. As shown in Fig. 3, an adaptive attention module consists of a frozen self-attention module and a learnable cross-attention module. Its self-attention weights are initialized from SD v1.5 to retain generative capacity. Given the query features $\mathbf{Z}_{n}$ from the Photography-Net and the animal identity features $\mathbf{F}_{a}$ from the Animal-Net, the output $\mathbf{O}_{h}$ of the adaptive attention module is defined as:
$\mathbf{O}_{h}=\underbrace{\operatorname{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}}_{\text{Self-attention}}+\lambda\underbrace{\operatorname{Softmax}\left(\frac{\mathbf{Q}\left(\mathbf{K}_{ID}\right)^{\top}}{\sqrt{d}}\right)\mathbf{V}_{ID}}_{\text{Cross-attention}}~,$ (1)
where $\lambda\in[0,1]$ controls the strength of the animal identity feature condition. $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are derived from $\mathbf{Z}_{n}$, while $\mathbf{K}_{ID}=\mathbf{F}_{a}\mathbf{W}_{k_{ID}}$ and $\mathbf{V}_{ID}=\mathbf{F}_{a}\mathbf{W}_{v_{ID}}$ are derived from $\mathbf{F}_{a}$. $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$, $\mathbf{W}_{k_{ID}}$, $\mathbf{W}_{v_{ID}}$ are the weights of the linear projection layers. To facilitate feature fusion, the self-attention and cross-attention share the query matrix $\mathbf{Q}$. The self-attention part is frozen, while the cross-attention part (i.e., $\mathbf{W}_{k_{ID}}$ and $\mathbf{W}_{v_{ID}}$) is trainable. This design effectively injects animal identity features while preserving the original T2I model’s [21] generative capabilities.
@@ 2.4 Frequency-Controlled Feature Integration Module
This module guides the diffusion process in the latent space through Discrete Cosine Transform (DCT) filtering (Fig. 3). First, a channel-wise 2D DCT is applied to the source domain latent features $\mathbf{L}_{0}$ (obtained from the VAE encoder) to acquire the frequency domain representation $\mathbf{F}_{DCT}$:
@@
Different frequency bands of the DCT spectrum encode distinct visual attributes. We design four types of DCT filters (masks) for mini-pass, low-pass, mid-pass, and high-pass filtering to manipulate visual properties from coarse structures to fine textures.
@@
$\qquad\times\cos\!\left(\frac{(2i+1)u\pi}{2h}\right)\cos\!\left(\frac{(2j+1)v\pi}{2w}\right).$
The $\mathbf{C}_{freq}$ signals obtained from mini-pass, low-pass, mid-pass, and high-pass filtering, respectively, control the texture, texture and structure, layout, and contour consistency between the generated and reference images.
@@ 2.5 Training and Objectives
AnimalBooth training is bifurcated into two distinct stages: First, the Animal-Net and Projection modules are trained. The objective function aims to minimize the Mean Squared Error between the predicted noise and the ground truth noise, while incorporating both text conditions $\mathbf{C}_{t}$ and animal identity features $\mathbf{C}_{a}$:
$L_{stage1}=\mathbb{E}_{\mathbf{z}_{t},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{C}_{t},\mathbf{C}_{a},t}\left\|\epsilon_{\theta}\left(\mathbf{z}_{t},\mathbf{C}_{t},\mathbf{C}_{a},t\right)-\epsilon_{t}\right\|^{2},$ (5)
@@
where $\mathbf{z}_{t}$ represents the noisy latent features at time step $t$, $\mathbf{C}_{t}$ is the text condition, and $\mathbf{C}_{freq}$ is computed by the frequency-controlled feature integration module based on the source domain latent features $\mathbf{L}_{0}$ (i.e., the VAE encoder output). The parameters of ControlNet are trained using four distinct DCT filters ($Mask_{mini},Mask_{low},Mask_{mid},Mask_{high}$).
During inference, classifier-free guidance [24] is employed to balance conditional and unconditional generation:
$\begin{split}\hat{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{C}_{t},\mathbf{C}_{a},\mathbf{C}_{freq},t)&=w\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{C}_{t},\mathbf{C}_{a},\mathbf{C}_{freq},t)\\ &+(1-w)\epsilon_{\theta}(\mathbf{x}_{t},t),\end{split}$ (7)
where $w$ is the guidance scale.
Fig. 4: Example results with different animal image strength $\lambda$.
@@ 3.1 Experimental Setup
We utilize Stable Diffusion v1.5 [21] as the pre-trained Latent Diffusion Model (LDM) and fine-tune it for personalized animal image generation. To comprehensively evaluate the model’s capabilities in generating high-definition animal images, we constructed a specialized AnimalBench dataset, comprising 10,958 training images and 1,000 test images. Each entry in the dataset consists of a high-definition, single-subject animal image–text pair, with examples depicted in Fig. 1. Training was conducted on a server equipped with eight NVIDIA A100-SXM4-80GB GPUs. We employed the AdamW optimizer with an initial learning rate of 1e-5 and a batch size of 4. The model was trained at a resolution of 512x512 pixels. During the inference phase, we used the DDIM sampler [25] for 50 sampling steps.

### v2

@@ Abstract
Personalized animal image generation is challenging due to rich appearance cues and large morphological variability. Existing approaches often exhibit feature misalignment across domains, which leads to identity drift. We present AnimalBooth, an inference-time tuning-free framework that strengthens identity preservation with an Animal-Net and an adaptive attention module, mitigating cross-domain alignment errors. We further introduce a frequency-controlled feature integration module that applies Discrete Cosine Transform filtering in the latent space to guide the diffusion process, enabling a coarse-to-fine progression from global structure to detailed texture. To advance research in this area, we curate AnimalBench, a high-resolution dataset for animal personalization. Extensive experiments show that AnimalBooth consistently outperforms strong baselines on multiple benchmarks with superior efficiency, improving both identity fidelity and perceptual quality. The code and dataset will be made publicly available in the future.
@@ 1 Introduction
Personalized multimodal generation is a prominent yet challenging subfield that aims to synthesize images conforming to both textual descriptions (text–image consistency) and the intrinsic characteristics of custom concepts (identity consistency) (Deng et al., 2025; Dai et al., 2025; Zhang et al., 2024b; Shen et al., 2025a). This paradigm shows strong potential across diverse applications ranging from creative artistry to product design (Shen et al., 2025d; c). However, accurately capturing the identity of animal subjects, which exhibit complex visual features such as fine-grained fur textures and non-rigid morphologies, remains a significant hurdle (Xu et al., 2025; Shen and Tang, 2024).
Animal personalization is uniquely challenging compared to human subject personalization due to several factors: (1) Non-rigid deformations: animals exhibit diverse poses and body configurations that vary dramatically across species; (2) Anatomical variations: skeletal structures and body proportions differ significantly between species (e.g., felines vs. bovines); (3) Intricate texture requirements: fur patterns, scales, and feathers require fine-grained preservation that general-purpose models often fail to maintain. As demonstrated in recent work (Xu et al., 2025), general-purpose models like DiT-based architectures frequently suffer from identity drift in these challenging scenarios.
@@
Existing personalization methodologies can be broadly categorized by their feature extraction and optimization strategies. The first category comprises optimization-based methods that achieve high fidelity by fine-tuning model components. A representative example is DreamBooth (Ruiz et al., 2022), which fine-tunes the UNet backbone, while Textual Inversion (Gal et al., 2022) instead optimizes word embeddings. This paradigm has been extended to support multiple customized concepts (Kumari et al., 2023; Ma et al., 2024; Wang et al., 2025), including applications such as virtual dressing and editing (Shen et al., 2025b). Despite their fidelity, such approaches demand substantial computational resources and often exhibit an inherent trade-off between fidelity and diversity due to overfitting. The second category, encoder-based and tuning-free methods, offers a more efficient alternative. For instance, IP-Adapter (Ye et al., 2023) and its successors (Zhang et al., 2024a) utilize frozen external encoders (e.g., CLIP) to inject identity features. While efficient, their performance is fundamentally limited by the representational capacity of the encoder, particularly when modeling animals with unique identity cues such as body structures and fur patterns. Consequently, identity distortion and detail loss are frequent. Other directions include retrieval-augmented generation (Chen et al., 2022) or LLM-enhanced frameworks (Zeng et al., 2024; Pan et al., 2023; Sun et al., 2024), yet the challenge of inadequate domain-specific representation remains.
Motivated by bridging this domain gap, we propose AnimalBooth, a feature-enhanced, inference-time tuning-free personalized generation framework specifically tailored for animals. To systematically address these identified challenges, our framework incorporates several targeted mechanisms: (1) To handle non-rigid deformations, we introduce a frequency-controlled feature integration module that leverages low-frequency signals to guide coarse structural consistency; (2) To overcome encoder capacity limits and identity drift, we design a dedicated Animal-Net with a Q-Former bottleneck to filter noise and capture subject-specific semantic features; (3) To maintain text-control while injecting identity, we employ a dual-path adaptive attention mechanism.
@@
Furthermore, existing datasets such as AFHQ (Choi et al., 2020) or MS-COCO are primarily designed for classification, translation, or general captioning, and lack the high-resolution subject masks and fine-grained captions necessary for precise animal personalization. To address this, we construct AnimalBench, a curated high-definition dataset for animal personalization. Our contributions are threefold: (1) We propose a dedicated branch architecture comprising a lightweight Animal-Net and a novel dual-path adaptive attention mechanism, effectively mitigating identity distortion while preserving generative capacity. (2) We design a frequency-controlled feature integration module leveraging DCT filtering in latent space to enhance attribute manipulation and texture fidelity. (3) We establish and release AnimalBench, a high-definition dataset for animal personalization, on which our method achieves state-of-the-art performance while being significantly more efficient than large-scale DiT-based models.
Figure 1: Qualitative comparison with SOTA methods.
@@
Figure 2: AnimalBooth primarily consists of a trainable Animal-Net and a frozen Photography-Net. The Animal-Net incorporates an Adaptive Attention module for efficient identity feature injection and a Frequency-Controlled module (based on ControlNet (Zhang et al., 2023)) for enhanced control over visual attributes like structure and texture, while the Photography-Net integrates these features with text prompts within the latent space.
Figure 3: Examples of AnimalBench dataset. Each instance provides five components: (a) a high-quality source image, (b) a detailed caption describing the scene and subject, (c) a precise pixel-level semantic segmentation mask of the primary subject, (d) the resulting masked subject for visual verification, and (e) the corresponding paired image.
@@ 2.1 Overall Architecture
As depicted in Fig. 2, AnimalBooth effectively integrates a trainable Animal-Net with a frozen Photography-Net. The Animal-Net directly refines the complex features and flexible morphologies of animals within the feature space, effectively bypassing the complexities of cross-domain alignment. Through adaptive attention, identity features are efficiently injected. The incorporation of a frequency-controlled module based on ControlNet (Zhang et al., 2023) further enhances fine-grained control over visual attributes such as the structure and textural details of the generated images. The frozen Photography-Net then merges these extracted fine-grained features with text prompts in the latent space.
@@ 2.2 Animal-Net
In the task of personalized animal image generation, precise extraction of fine-grained animal identity features is paramount for maintaining consistency between generated images and reference animals. To achieve this, we propose a specialized Animal-Net capable of simultaneously capturing both semantic information and high-frequency textural features of animals. Specifically, given an animal reference image $\mathcal{X}_{a}\in\mathbb{R}^{3\times H\times W}$, we first transform it into a latent space representation $\mathbf{L}_{a}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ using a frozen VAE encoder (Rombach et al., 2022). We then extract token embeddings from $\mathcal{X}_{a}$ using a frozen CLIP image encoder (Radford et al., 2021) and a trainable projection layer.
Why Q-Former? We employ Q-Former (Li et al., 2023b) as the projection layer due to its superior ability to act as a semantic bottleneck that effectively filters background noise while capturing intricate animal textures via dynamic attention. Unlike static mapping approaches (e.g., feature concatenation or MLP adapters), Q-Former’s learnable query tokens dynamically attend to relevant identity features. As shown in Table 2, Q-Former significantly outperforms simpler alternatives in identity fidelity metrics.
Subsequently, animal features within the Animal-Net interact extensively through a cross-attention mechanism, similar to the interaction between text and image features in the original Text-to-Image (T2I) model (Rombach et al., 2022). This interaction ensures a deep fusion of semantic and textural features. Finally, the output of the Animal-Net is aligned in parallel with the Photography-Net, injecting these fine-grained animal identity features into the Photography-Net via an adaptive attention module. Notably, the Animal-Net is solely used for encoding reference images; therefore, during the diffusion process, no noise is added to the reference image, and it undergoes only a single forward pass.
@@ 2.3 Adaptive Attention Module
To enable personalized animal image generation, the Photography-Net must possess both its original generative capabilities and the ability to fuse animal identity features. We freeze the core modules of the Photography-Net to preserve the former, and achieve the latter through an adaptive attention module. The architecture of the Photography-Net in AnimalBooth builds upon SD v1.5 (Rombach et al., 2022), with all self-attention modules replaced by adaptive attention modules.
Dual-Path Architecture. Unlike standard single-path cross-attention strategies commonly used in subject-driven generation, our adaptive attention module employs a novel dual-path architecture: a frozen self-attention path (to preserve the original generative capacity of Stable Diffusion) and a trainable cross-attention path (to inject identity features), both sharing a single Query matrix $\mathbf{Q}$. This shared-query design enables seamless feature fusion while maintaining the pre-trained model’s semantic understanding. As shown in Fig. 2, an adaptive attention module consists of a frozen self-attention module and a learnable cross-attention module. Its self-attention weights are initialized from SD v1.5 to retain generative capacity. Given the query features $\mathbf{Z}_{n}$ from the Photography-Net and the animal identity features $\mathbf{F}_{a}$ from the Animal-Net, the output $\mathbf{O}_{h}$ of the adaptive attention module is defined as:
$\mathbf{O}_{h}=\underbrace{\operatorname{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}}_{\text{Frozen Self-attention}}+\lambda\underbrace{\operatorname{Softmax}\left(\frac{\mathbf{Q}\left(\mathbf{K}_{ID}\right)^{\top}}{\sqrt{d}}\right)\mathbf{V}_{ID}}_{\text{Trainable Cross-attention}}~,$ (1)
where $\lambda\in[0,1]$ is a dynamic coefficient that controls the strength of the animal identity feature condition, enabling flexible balance between identity preservation and creative generation. $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are derived from $\mathbf{Z}_{n}$, while $\mathbf{K}_{ID}=\mathbf{F}_{a}\mathbf{W}_{k_{ID}}$ and $\mathbf{V}_{ID}=\mathbf{F}_{a}\mathbf{W}_{v_{ID}}$ are derived from $\mathbf{F}_{a}$. The self-attention part is frozen, while the cross-attention part (i.e., $\mathbf{W}_{k_{ID}}$ and $\mathbf{W}_{v_{ID}}$) is trainable. This design effectively injects animal identity features while preserving the original T2I model’s (Rombach et al., 2022) generative capabilities.
@@ 2.4 Frequency-Controlled Feature Integration Module
This module guides the diffusion process in the latent space through Discrete Cosine Transform (DCT) filtering, implemented via ControlNet (Zhang et al., 2023) (Fig. 2). First, a channel-wise 2D DCT is applied to the source domain latent features $\mathbf{L}_{0}$ (obtained from the VAE encoder) to acquire the frequency domain representation $\mathbf{F}_{DCT}$:
@@
Different frequency bands of the DCT spectrum encode distinct visual attributes: low-frequency components capture global structure and identity-related morphology, while high-frequency components encode fine textures and edges. We design four types of DCT filters (masks) for mini-pass, low-pass, mid-pass, and high-pass filtering to manipulate visual properties from coarse structures to fine textures.
@@
$\quad\times\cos\!\left(\frac{(2i+1)u\pi}{2h}\right)\cos\!\left(\frac{(2j+1)v\pi}{2w}\right).$
The $\mathbf{C}_{freq}$ signals obtained from mini-pass, low-pass, mid-pass, and high-pass filtering, respectively, control the texture, texture and structure, layout, and contour consistency between the generated and reference images. As our experiments demonstrate (Table 4), low-pass filtering achieves optimal identity preservation by retaining the structure and morphology information critical for animal recognition.
@@ 2.5 Training and Objectives
AnimalBooth training is bifurcated into two complementary stages that follow a coarse-to-fine progression:
Stage 1: Identity Learning. The Animal-Net and Projection modules are trained to capture global subject identity. The objective function aims to minimize the Mean Squared Error between the predicted noise and the ground truth noise, while incorporating both text conditions $\mathbf{C}_{t}$ and animal identity features $\mathbf{C}_{a}$:
$L_{stage1}=\mathbb{E}_{\mathbf{z}_{t},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{C}_{t},\mathbf{C}_{a},t}\left\|\epsilon_{\theta}\left(\mathbf{z}_{t},\mathbf{C}_{t},\mathbf{C}_{a},t\right)-\epsilon_{t}\right\|^{2}.$ (5)
Stage 2: Texture Enhancement. The frequency-controlled ControlNet module is trained to enhance fine textures and structural details. This decoupled training strategy avoids convergence difficulties that arise from jointly optimizing identity and texture objectives:
@@
where $\mathbf{z}_{t}$ represents the noisy latent features at time step $t$, $\mathbf{C}_{t}$ is the text condition, and $\mathbf{C}_{freq}$ is computed by the frequency-controlled feature integration module based on the source domain latent features $\mathbf{L}_{0}$. The parameters of ControlNet are trained using four distinct DCT filters ($Mask_{mini},Mask_{low},Mask_{mid},Mask_{high}$).
During inference, classifier-free guidance (Ho and Salimans, ) is employed to balance conditional and unconditional generation:
$\begin{split}\hat{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{C}_{t},\mathbf{C}_{a},\mathbf{C}_{freq},t)&=w\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{C}_{t},\mathbf{C}_{a},\mathbf{C}_{freq},t)+(1-w)\epsilon_{\theta}(\mathbf{x}_{t},t),\end{split}$ (7)
where $w$ is the guidance scale. We empirically set $w=7.5$ in our experiments (see Table 2 for ablation).
Figure 4: Example results with different animal image strength $\lambda$.
@@ 3.1 Experimental Setup
We utilize Stable Diffusion v1.5 (Rombach et al., 2022) as the pre-trained Latent Diffusion Model (LDM) and train it for personalized animal image generation. To comprehensively evaluate the model’s capabilities in generating high-definition animal images, we constructed a specialized AnimalBench dataset, comprising 10,958 training images and 1,000 test images. Each entry in the dataset consists of a high-definition, single-subject animal image–text pair, with examples depicted in Fig. 3. Training was conducted on a server equipped with eight NVIDIA A100-SXM4-80GB GPUs. We employed the AdamW optimizer with an initial learning rate of 1e-5 and a batch size of 4. The model was trained at a resolution of $\times512$ pixels. During the inference phase, we used the DDIM sampler (Ho et al., 2020) for 50 sampling steps with guidance scale $w=7.5$.